OpenStudy (anonymous):

The length of a side of an equilateral triangle is the same as the length of a rectangle and the width of the rectangle is 2 inches less than its length, if the perimeter of the triangle s 4 inches less than the perimeter of the rectangle, what are the dimensions of the rectangle?

OpenStudy (anonymous):

Help please!?

Parth (parthkohli):

Let one side in the equilateral triangle be \(x\). As we know from the definition of perimeter, it's the sum of all sides. We do know that all sides are \(x\) because it's an equilateral triangle.

Parth (parthkohli):

The perimeter of that equilateral triangle is \(x + x + x\) or \(3x\), right?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Okay, well, now let's go to the rectangle. We know that the length of the rectangle is the same as the side in our equilateral triangle. The length becomes \(x\) then, right?

OpenStudy (anonymous):

yes

Parth (parthkohli):

OK, now we know that the width is 2 inches less than the length, so width becomes \(x - 2\). Okay?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Do you know what the formula is for finding out the perimeter of a rectangle?

OpenStudy (anonymous):

2(l+w)

Parth (parthkohli):

That's right! Now, we have to put in \(x\) for length and \(x - 2\) for width.\[\implies 2(x + x - 2) = Perimeter \]

OpenStudy (anonymous):

So, 2x+2x+4?

Parth (parthkohli):

Wait! Let's just keep that like it is.\[2(x + x - 2) = 3x - 2 \]Agree with this?

OpenStudy (anonymous):

oh ok. yes

Parth (parthkohli):

Actually, it'd be\[ 2(x + x - 2) - 4 = 3x\]

Parth (parthkohli):

Now solve the equation.

OpenStudy (anonymous):

2x+2x-4-4=3x 4x+8=3x is that right so far?

Parth (parthkohli):

No... :I

OpenStudy (anonymous):

No? how?

Parth (parthkohli):

\[2(x + x - 2) - 4 = 3x\]Maybe simplify the parentheses.\[2(2x - 2) - 4 = 3x\]\[ 4x - 4 - 4 = 3x\]\[ 4x - 8 = 3x\]

Parth (parthkohli):

As to your question, no, that's not the answer because when you complete solving that equation, you get \(x = -8\) which is not possible. Why? There cannot be a negative side! Remember how \(x = \) side of the triangle?

OpenStudy (anonymous):

oh ok. whats next?

Parth (parthkohli):

You get \(x = 8\).

Parth (parthkohli):

\(x = \text{length of the rectangle }=8\\ x -2 = \text{width of the rectangle} = 8 - 2 = 6\)

OpenStudy (anonymous):

wait can we go back to 4x-8=3x. how did you get from there to x=8?

Parth (parthkohli):

\[\begin{array}{lr} 4x - 8 &=&3x\\ 4x - 8 + 8&=&3x + 8\\ 4x &=&3x + 8\\4x - 3x&=&3x + 8 - 3x\\ x &=&8 \end{array}\]

OpenStudy (anonymous):

oh ok. thats what i thought just checking.

OpenStudy (anonymous):

whats next?

Parth (parthkohli):

I wrote it above.

OpenStudy (anonymous):

so the width of the rectangle is 6 and the length is 8?